Does anyone understand this question?

The standard deviation of a distribution is 20. If a sample of 225 is selected, what is the standard error of the mean?

Answer by stanbon(48568)

The standard deviation of a distribution is 20. If a sample of 225 is selected, what is the standard error of the mean?
--------------------
20/sqrt(225) = 20/15 = 4/3 = 1.33333333....
===============
Cheers,
Stan H.